# Matryoshka Quantization

### v1

@@ Abstract
Quantizing model weights is critical for reducing the communication and inference costs of large models. However, quantizing models – especially to low precisions like int4 or int2 – requires a trade-off in model quality; int2, in particular, is known to severely degrade model quality. Consequently, practitioners are often forced to maintain multiple models with different quantization levels or serve a single model that best satisfies the quality-latency trade-off. On the other hand, integer data types, such as int8, inherently possess a nested (Matryoshka) structure where smaller bit-width integers, like int4 or int2, are nested within the most significant bits. This paper proposes Matryoshka Quantization (${\rm MatQuant}$), a novel multi-scale quantization technique that addresses the challenge of needing multiple quantized models. It allows training and maintaining just one model, which can then be served at different precision levels. Furthermore, due to the co-training and co-distillation regularization provided by ${\rm MatQuant}$, the int2 precision models extracted by ${\rm MatQuant}$ can be up to $10\%$ more accurate than standard int2 quantization (using techniques like QAT or OmniQuant). This represents significant progress in model quantization, demonstrated by the fact that, with the same recipe, an int2 FFN-quantized Gemma-2 9B model is more accurate than an int8 FFN-quantized Gemma-2 2B model.
@@ 1 Introduction
Quantizing weights and/or activations can significantly reduce the overall communication load and is, therefore, one of the most popular techniques for reducing inference costs (Dettmers et al., 2022). While floating-point representations are standard for training, integer data types such as int8, int4, and int2 are appealing alternatives for inference. However, current methods for quantizing to these varying integer precisions typically treat each target precision as an independent optimization problem, leading to a collection of distinct models rather than a single, versatile one. Furthermore, quantizing to extremely low precisions like int2 is known to be highly inaccurate. In this work, we pose the question of whether both of the above challenges can be addressed; that is, can we train a single model from which we can extract multiple accurate lower-precision models? We answer this question in the affirmative by introducing Matryoshka Quantization (${\rm MatQuant}$), a novel multi-scale training method that leverages the inherent nested (Matryoshka) structure (Kusupati et al., 2022) within integer data types (Figure 1a). Specifically, slicing the top bits of an int8-quantized weight can directly yield an int4 or int2 model. Existing quantization techniques often neglect this structure, which limits the potential for multi-scale adaptable models operating at various bit-widths with optimal performance.
@@
Instead, ${\rm MatQuant}$ simultaneously optimizes model weights across multiple precision levels (e.g., int8, int4, int2). At a high level, we represent each model parameter at different precision levels using shared most significant bits (MSBs), and then jointly optimize the loss for each precision level. This allows us to develop a single quantized model that can effectively operate at any of the chosen bit-widths, offering a spectrum of accuracy-versus-cost options. ${\rm MatQuant}$ is a general-purpose technique, applicable to most learning-based quantization methods, such as Quantization Aware Training (QAT) (Jacob et al., 2018) and OmniQuant (Shao et al., 2023).
@@
We demonstrate the efficacy of ${\rm MatQuant}$ when applied to quantizing the Feed-Forward Network (FFN) parameters of standard LLMs (Gemma-2 2B, 9B, and Mistral 7B) (Vaswani et al., 2017) – typically, FFN is the main latency block hence the focus on improving the most significant component’s latency. Our results show that ${\rm MatQuant}$ produces int8 and int4 models with comparable accuracy to independently trained baselines, despite the benefit of shared model parameters. Critically, the int2 models generated by ${\rm MatQuant}$ significantly outperform their individually trained counterparts, with 8% higher accuracy on downstream tasks (Figure 1b). We also extend ${\rm MatQuant}$ to quantize all weights of a Transformer layer. Finally, we find that quantizing with ${\rm MatQuant}$ shifts the quantized weight distribution toward higher values, contributing to improved int2 performance (Figure1c).
@@
Beyond improving chosen precision performance, ${\rm MatQuant}$ allows for seamless extraction of interpolative bit-widths, such as int6 and int3. ${\rm MatQuant}$ also admits a dense accuracy-vs-cost pareto-optimal trade-off by enabling layer-wise Mix’n’Match of different precisions. This ensures deployment of say an effective int3 sized model even if the underlying hardware only supports int4 and int2. Overall, ${\rm MatQuant}$ and its variants present a significant step toward developing multi-scale models with high flexibility and performance, pushing the boundaries of low-bit quantization for efficient LLM inference.
@@ 2 Related Work
Training across multiple data scales (resolutions) was heavily popularized in computer vision for both recognition and generation (Adelson et al., 1984; Lin et al., 2017; Denton et al., 2015). More recently, the paradigm of multi-scale training has shifted to models (Rippel et al., 2014; Yu et al., 2018; Kusupati et al., 2022; Devvrit et al., 2023), where the data remains the same, and models of varying capacity, all nested within one large model, are trained jointly. This joint, nested (Matryoshka-style) learning with varying model sizes results in a smooth accuracy-vs-compute trade-off and is beneficial in many downstream applications and real-world deployments. However, the most obvious structure with a nested nature is the bit structure of the integer data type. Given the success of multi-scale training for inputs, outputs, and model weights, it is imperative to explore it further for integer data types, especially in the context of quantization, which aids in the deployment of resource-intensive LLMs.
@@ 3.2 ${\rm MatQuant}$
Once we have this structure, we can optimize for several precisions by slicing the MSBs from the largest bit-width we are optimizing for. Let $R=\{r_{1},r_{2},...,r_{K}\}$ be the bit-widths we want to optimize for, $Q(\cdot,)$ represent the quantization function of the base algorithm (i.e., any learning-based quantization scheme), $\mathcal{L}(\cdot)$ represent the loss function pertaining to the base algorithm, $F(\cdot)$ represent the forward pass required to compute the loss, $\theta$ represent the set of model/auxiliary parameters we are optimizing for and let $W_{F}$ represent the model parameters. ${\rm MatQuant}$’s overall objective can be formulated as follows:
@@
In this work, we default to training ${\rm MatQuant}$ with three bit-widths, $R=\{8,4,2\}$, and subsequently perform a grid search over $\lambda_{r}$. This process aims to optimize performance such that the model performs well across all targeted precision levels. Further, while the focus of this paper is primarily on integer data types, we discuss the possibility of extending ${\rm MatQuant}$ to floating-point representations in Section 5.5.
@@ 3.2.1 Interpolative Behavior
${\rm MatQuant}$ also enables the use of different precisions at different layers through layer-wise Mix’n’Match (Devvrit et al., 2023), even though we never trained for these combinatorial possibilities. These large number of models, obtained at no cost, densely span the accuracy-vs-memory trade-off. We explore several Mix’n’Match strategies and find that having a higher precision (int8) in the middle layers and a lower precision (int2) at the start and end is Pareto-optimal among hundreds of possible models. See Section 4.3 for detailed experiments.
@@ 4 Experiments
In this section, we present an empirical evaluation of ${\rm MatQuant}$ working with two popular learning-based quantization methods: OmniQuant (Section 4.1) and QAT (Section 4.2). We demonstrate ${\rm MatQuant}$’s efficiency on Transformer-based LLMs. Unless otherwise mentioned, our primary focus is on weight quantization within the parameter-intensive FFN blocks of the Transformer layer.
@@ 4.1 ${\rm MatQuant}$ with OmniQuant
Table 1 shows the efficacy of ${\rm MatQuant}$ when used with FFN-only OmniQuant and compared to explicitly trained OmniQuant baselines for the target precisions, i.e., int8, int4, and int2, across all the models. While the average downstream accuracy of ${\rm MatQuant}$ for int8 and int4 quantization is within 0.5% of the corresponding independently trained baselines, the int2 quantized models of ${\rm MatQuant}$ are $4.37\%$, $8.01\%$, and $6.35\%$ more accurate for Gemma-2 2B, 9B, and Mistral 7B, respectively. Similar trends and improvements follow when measuring performance through validation log perplexity. Further, the quantized int4 and int2 models sliced from the int8 OmniQuant baseline suffer a significant drop in accuracy around int4, demonstrating that the nested structure of int8 is not well utilized.
@@
Beyond the target quantization granularities (int8, int4, and int2), ${\rm MatQuant}$ allows for bit-width interpolation to bit-widths not optimized during training. We find that the accuracy of the int6 and int3 models obtained by slicing the ${\rm MatQuant}$ models is comparable to explicitly trained baselines for both precisions.
@@ 4.2 ${\rm MatQuant}$ with QAT
To further demonstrate the generality of ${\rm MatQuant}$, we experiment on the same models using the popular QAT technique. Following the trend of experimental results with OmniQuant, we show in Table 2 that the models trained using ${\rm MatQuant}$ with QAT are comparable to the explicitly trained baselines for all the targeted bit-widths of int8 and int4. However, int2 quantized models using ${\rm MatQuant}$ are $4.69\%$, $6.30\%$, and $6.34\%$ more accurate for Gemma-2 2B, 9B, and Mistral 7B, respectively.
@@
Figure 2: Mix’n’Match on Gemma-2 9B model trained using ${\rm MatQuant}$ with OmniQuant allows elastic pareto-optimal accuracy-vs-cost model extraction for free during deployment.
@@ 4.3 Layerwise Mix’n’Match
Alongside the strong slicing-based interpolative properties, quantization with ${\rm MatQuant}$ also enables another form of elastic and interpolative behavior through Mix’n’Match. Mix’n’Match provides a mechanism to obtain a combinatorial number of strong models by using different quantization granularities, from the target bit-widths – i.e., int8, int4, and int2 across layers. Figure 2 shows the ability of Mix’n’Match to densely span the Pareto-optimal accuracy-vs-bits-per-FFN-parameter (memory/cost) trade-off for the Gemma-2 9B model trained using ${\rm MatQuant}$ with OmniQuant – sometimes even improving on the bfloat16 model accuracy. While there are many more feasible models, we only showcase the best models obtained through the strategy described in Section 3.2.1 and further expanded in Appendix A. Interestingly, the Mix’n’Match models with effective bit-width of 3 and 6 are as accurate as models obtained through slicing. This opens up possibilities for effective serving depending on hardware support (Section 5.4).
@@ 5 Ablations and Discussion
int8 $(1,1,1)$ $67.42$ $73.97$ $73.46$
$(1,1,\sqrt{2})$ $67.31$ $73.45$ $73.41$
$(2,2,1)$ $\bf 67.85$ $74.02$ $\bf 73.82$
$(\sqrt{2},\sqrt{2},1)$ $67.3$ $\bf 74.33$ $\bf 73.82$
int4 $(1,1,1)$ $66.11$ $73.88$ $73.13$
$(1,1,\sqrt{2})$ $\bf 66.70$ $73.75$ $73.29$
$(2,2,1)$ $66.54$ $\bf 74.33$ $\bf 73.5$
$(\sqrt{2},\sqrt{2},1)$ $66.46$ $74.26$ $72.97$
int2 $(1,1,1)$ $\bf 55.71$ $\bf 68.52$ $65.99$
$(1,1,\sqrt{2})$ $57.08$ $67.93$ $\bf 66.28$
$(2,2,1)$ $\bf 55.70$ $66.72$ $63.49$
$(\sqrt{2},\sqrt{2},1)$ $55.29$ $68.25$ $57.85$
@@ 5.1 Weightings ($\lambda_{r}$) for ${\rm MatQuant}$
Depending on the constraints, we may wish to maximize the accuracy of one of the target bit-widths in ${\rm MatQuant}$. Equation 7 provides a general formulation of ${\rm MatQuant}$ that supports a grid search on the weights $\lambda_{r}$ for bit-width $r$. The results in Section 4 are with the weights that have balanced performance across target precisions. Table 3 shows the weight multiplier ablation results for Gemma-2 2B, 9B, and Mistral 7B. While equal weighting for all precisions works well, we see that higher weights for a specific precision results in increased accuracy for that bit-width. This re-weighting to improve int8 and int4 models often results in a minor accuracy drop for the int2 models. We can consider re-weighting as scaling the importance of the bits during training, and finding an optimal grid-search-free recipe is an interesting research question.

### v2

@@ Abstract
Quantizing model weights is critical for reducing the communication and inference costs of large models. However, quantizing models – especially to low precisions like int4 or int2 – requires a trade-off in model quality; int2, in particular, is known to severely degrade model quality. Consequently, practitioners are often forced to maintain multiple models with different quantization levels or serve a single model that best satisfies the quality-latency trade-off. On the other hand, integer data types, such as int8, inherently possess a nested (Matryoshka) structure where smaller bit-width integers, like int4 or int2, are nested within the most significant bits. Leveraging this insight, in this paper, we propose Matryoshka Quantization (${\rm MatQuant}$), a novel multi-scale quantization technique that alleviates the aforementioned challenge. This technique allows us to train and maintain a single quantized model but serve it with the precision demanded by the deployment. Furthermore, leveraging ${\rm MatQuant}$’s co-training and co-distillation regularization, int2 precision models extracted by ${\rm MatQuant}$ outperform standard int2 quantization by up to to 4% and 7% with OmniQuant and QAT as base algorithms respectively. Finally, we demonstrate that by using an extra bit to represent outliers, a model with an effective precision of 2.05-bit gives an additional 6% improvement with OmniQuant as the base algorithm.
@@ 1 Introduction
Quantizing weights and/or activations can significantly reduce the overall communication load and is, therefore, one of the most popular techniques for reducing inference costs (Dettmers et al., 2022). While floating-point representations are standard for training, integer data types such as int8, int4, and int2 are appealing alternatives for inference. However, current methods for quantizing to these varying integer precisions typically treat each target precision as an independent optimization problem, leading to a collection of distinct models rather than a single, versatile one. Furthermore, quantizing to extremely low precisions like int2 is known to be highly inaccurate. In this work, we pose the question of whether both of the above challenges can be addressed; that is, can we train a single model from which we can extract multiple accurate lower-precision models? We answer this question in the affirmative by introducing Matryoshka Quantization (${\rm MatQuant}$), a novel multi-scale training method that leverages the inherent nested (Matryoshka) structure (Kusupati et al., 2022) within integer data types (Figure 1a). Specifically, slicing the most significant bits (MSBs) of an int8-quantized weight can directly yield an int4 or int2 model. Existing quantization techniques often neglect this structure, which limits the potential for multi-scale adaptable models operating at various bit-widths with optimal performance.
@@
Instead, ${\rm MatQuant}$ simultaneously optimizes model weights across multiple precision levels (e.g., int8, int4, int2). At a high level, we represent each model parameter at different precision levels using shared MSBs, and then jointly optimize the loss for each precision level. This allows us to develop a single quantized model that can effectively operate at any of the chosen bit-widths, offering a spectrum of accuracy-vs-cost options. ${\rm MatQuant}$ is a general-purpose technique, applicable to most learning-based quantization methods, such as Quantization Aware Training (QAT) (Jacob et al., 2018) and OmniQuant (Shao et al., 2023).
@@
We demonstrate the efficacy of ${\rm MatQuant}$ when applied to quantizing the Feed-Forward Network (FFN) parameters of standard LLMs (Gemma-2 2B, 9B, and Mistral 7B) (Vaswani et al., 2017) – typically, FFN is the main latency block hence the focus on improving the most significant component’s latency. Our results show that ${\rm MatQuant}$ produces int8 and int4 models with comparable accuracy to independently trained baselines, despite the benefit of shared model parameters. Critically, the int2 models generated by ${\rm MatQuant}$ significantly outperform their individually trained counterparts, with $4$% higher accuracy on downstream tasks (Figure 1b). We also extend ${\rm MatQuant}$ to quantize all weights of a Transformer layer. In Figure 1c, we find that quantizing with ${\rm MatQuant}$ shifts the quantized weight distribution toward higher values, contributing to improved int2 performance. Finally, in Section 7, we also demonstrate that using an extra bit to represent outliers significantly boosts the performance for our sliced int2 models.
@@
Beyond improving chosen precision performance, ${\rm MatQuant}$ allows for seamless extraction of interpolative bit-widths, such as int6 and int3. ${\rm MatQuant}$ also admits a dense accuracy-vs-cost trade-off by enabling layer-wise Mix’n’Match of different precisions. Therefore, even if the hardware only supports int4 and int2, it’s possible to serve models at various effective precisions, tailored to the deployment environment. Overall, ${\rm MatQuant}$ and its variants present a significant step toward developing multi-scale models with high flexibility and performance, pushing the boundaries of low-bit quantization for efficient LLM inference.
@@ 2 Related Work
Training across multiple data scales (resolutions) was heavily popularized in computer vision for both recognition and generation (Adelson et al., 1984; Lin et al., 2017; Denton et al., 2015). More recently, the paradigm of multi-scale training has shifted to models (Rippel et al., 2014; Yu et al., 2018; Kusupati et al., 2022; Devvrit et al., 2023), where the data remains the same, and models of varying capacity, all nested within one large model, are trained jointly. This joint, nested (Matryoshka-style) learning with varying model sizes results in a smooth accuracy-vs-compute trade-off and is beneficial in many downstream applications and real-world deployments. However, the most obvious structure with a nested nature is the bit structure of the integer data type. Given the success of multi-scale training for inputs, outputs, and model weights, it is imperative to explore it further for integer data types, especially in the context of quantization, which aids in the deployment of resource-intensive LLMs. Following this idea, Yu et al. (2019) have successfully trained a single model that can do well at any precision. However, the experiments were limited to ConvNets and small Neural Networks. In this paper, we extend the idea of nested precision to LLMs and show that it indeed works at scale. We also show that, for the first time, our models are quality neutral for intermediate precisions such as int3 and int6 that we never trained for, and densely span the accuracy-vs-bits trade-off. In Section 5.3, we show that even to train models for a fixed target precision, having loss over the sliced bits of an 8-bit model does better than training a model explicitly for that precision, indicating that ${\rm MatQuant}$ is a fundamentally better way to do low-bit quantization.
@@ 3.2 ${\rm MatQuant}$
$S(q^{c},r)=\text{clamp}\left(\left\lfloor\frac{q^{c}}{2^{c-r}}\right\rceil,0,2^{r}-1\right)*2^{c-r}\vspace{-3mm}$ (6)
Note that $\text{clamp}(\cdot)$ is required to curtail overflows generated by rounding. More details can be found in Appdendix A. Once we have this structure, we can optimize for several precisions by slicing the MSBs from the largest bit-width we are optimizing for. Let $R=\{r_{1},r_{2},...,r_{K}\}$ be the bit-widths we want to optimize for, $Q(\cdot,)$ represent the quantization function of the base algorithm (i.e., any learning-based quantization scheme), $\mathcal{L}(\cdot)$ represent the loss function pertaining to the base algorithm, $F(\cdot)$ represent the forward pass required to compute the loss, $\theta$ represent the set of model/auxiliary parameters we are optimizing for and let $W_{F}$ represent the model parameters. ${\rm MatQuant}$’s overall objective can be formulated as follows:
@@
In this work, we default to training ${\rm MatQuant}$ with three bit-widths, $R=\{8,4,2\}$, and subsequently perform a linear search over $\lambda_{r}$. This process aims to optimize performance such that the model performs well across all targeted precision levels. Further, while the focus of this paper is primarily on integer data types, we discuss the possibility of extending ${\rm MatQuant}$ to floating-point representations in Section 5.5.
@@ 3.2.1 Interpolative Behavior
${\rm MatQuant}$ also enables the use of different precisions at different layers through layer-wise Mix’n’Match (Devvrit et al., 2023), even though we never trained for these combinatorial possibilities. These large number of models, obtained at no cost, densely span the accuracy-vs-memory trade-off. We explore several Mix’n’Match strategies and find that having a higher precision (int8) in the middle layers and a lower precision (int2) at the start and end is the most optimal among hundreds of possible models. See Section 4.3 for detailed experiments.
@@ 4 Experiments
In this section, we present an empirical evaluation of ${\rm MatQuant}$ working with two popular learning-based quantization methods: OmniQuant (Section 4.1) and QAT (Section 4.2). We demonstrate ${\rm MatQuant}$’s efficiency on Transformer-based LLMs. Unless otherwise mentioned, our primary focus is on weight only quantization within the parameter-intensive FFN blocks of the Transformer layer.
@@ 4.1 ${\rm MatQuant}$ with OmniQuant
Table 1 shows the efficacy of ${\rm MatQuant}$ when used with FFN-only OmniQuant and compared to explicitly trained OmniQuant baselines for the target precisions, i.e., int8, int4, and int2, across all the models. While the average downstream accuracy of ${\rm MatQuant}$ for int8 and int4 quantization is within 0.5% of the corresponding independently trained baselines, the int2 quantized models of ${\rm MatQuant}$ are $1.04\%$, $3.11\%$, and $3.01\%$ more accurate for Gemma-2 2B, 9B, and Mistral 7B, respectively. Similar trends and improvements follow when measuring performance through validation log perplexity. Further, the quantized int4 and int2 models sliced from the int8 OmniQuant baseline suffer a significant drop in accuracy around int4, demonstrating that the nested structure of int8 is not well utilized.
@@
Beyond the target quantization granularities (int8, int4, and int2), ${\rm MatQuant}$ allows for bit-width interpolation to bit-widths not optimized during training. We find that the accuracy of the int6 and int3 models obtained by slicing the ${\rm MatQuant}$ models is comparable to their explicitly trained baselines.
@@ 4.2 ${\rm MatQuant}$ with QAT
To further demonstrate the generality of ${\rm MatQuant}$, we experiment on the same models using the popular QAT technique. Following the trend of experimental results with OmniQuant, we show in Table 2 that the models trained using ${\rm MatQuant}$ with QAT are comparable to the explicitly trained baselines for all the targeted bit-widths of int8 and int4. However, int2 quantized models using ${\rm MatQuant}$ are $4.46\%$, $6.27\%$, and $7.02\%$ more accurate for Gemma-2 2B, 9B, and Mistral 7B, respectively.
@@
Figure 2: Mix’n’Match on Gemma-2 9B model trained using ${\rm MatQuant}$ with OmniQuant allows elastic accuracy-vs-cost model extraction for free during deployment.
@@ 4.3 Layerwise Mix’n’Match
Alongside the strong slicing-based interpolative properties, quantization with ${\rm MatQuant}$ also enables another form of elastic and interpolative behavior through Mix’n’Match. Mix’n’Match provides a mechanism to obtain a combinatorial number of strong models by using different quantization granularities, from the target bit-widths – i.e., int8, int4, and int2 across layers. Figure 2 shows the ability of Mix’n’Match to densely span the accuracy-vs-bits-per-FFN-parameter (memory/cost) trade-off for the Gemma-2 9B model trained using ${\rm MatQuant}$ with OmniQuant. While there are many more feasible models, we only showcase the best models obtained through the strategy described in Section 3.2.1 and further expanded in Appendix B. Interestingly, the Mix’n’Match model, with a sub-4-bit effective width, is more accurate than the 4-bit sliced model. This opens up possibilities for effective serving depending on hardware support. Section 5.4 continues this discussion in greater depth.
@@ 5 Ablations and Discussion
int8 $(0.1,0.1,1)$ $\bf 68.02$ $\bf 74.05$ $73.27$
$(0.2,0.2,1)$ $67.91$ $73.91$ $73.44$
$(0.3,0.3,1)$ $68.01$ $73.88$ $73.56$
$(0.4,0.4,1)$ $67.95$ $73.84$ $\bf 73.65$
int4 $(0.1,0.1,1)$ $66.58$ $73.83$ $72.76$
$(0.2,0.2,1)$ $67.47$ $73.8$ $73.16$
$(0.3,0.3,1)$ $66.97$ $73.25$ $73.47$
$(0.4,0.4,1)$ $\bf 67.48$ $\bf 74.32$ $\bf 73.66$
int2 $(0.1,0.1,1)$ $\bf 52.37$ $63.35$ $63.25$
$(0.2,0.2,1)$ $51.88$ $64.04$ $\bf 63.99$
$(0.3,0.3,1)$ $51.05$ $\bf 64.1$ $63.6$
$(0.4,0.4,1)$ $51.69$ $61.98$ $62.75$
@@ 5.1 Weightings ($\lambda_{r}$) for ${\rm MatQuant}$
Depending on the constraints, we may wish to maximize the accuracy of one of the target bit-widths in ${\rm MatQuant}$. Equation 7 provides a general formulation of ${\rm MatQuant}$ that supports searching over the weight $\lambda_{r}$ for bit-width $r$. The results in Section 4 are with the weights that have balanced performance across target precisions. Table 3 shows the weight multiplier ablation results for Gemma-2 2B, 9B, and Mistral 7B. We find that a higher relative value for $\lambda_{2}$ is essential in attaining good int2 performance. Increasing $\lambda_{4},\lambda_{8}$ to improve int8 and int4 models often results in accuracy drop for the int2 models. In general, we can see that a higher relative weight for a specific precision results in increased accuracy for that bit-width. We can consider re-weighting as scaling the importance of the bits during training, and finding an optimal re-weighting recipe is an interesting research question.
[…]
Gemma-2 9B OmniQuant QAT
Data type Config. Task Avg. log pplx. Task Avg. log pplx.
int8 $[8,4,2]$ $\bf 74.05$ $2.438$ $74.52$ $2.262$
$[8,4,8\rightarrow 2]$ $72.76$ $2.473$ $74.75$ $2.242$
$[8,4,2,8\rightarrow 2]$ $73.99$ $\bf 2.435$ $\bf 74.87$ $\bf 2.240$
$[8,4,2,8\rightarrow 4;2]$ $73.85$ $2.437$ $74.81$ $2.240$
int4 $[8,4,2]$ $\bf 73.83$ $2.491$ $73.24$ $2.295$
$[8,4,8\rightarrow 2]$ $72.65$ $2.519$ $73.76$ $2.279$
$[8,4,2,8\rightarrow 2]$ $73.63$ $2.486$ $73.77$ $\bf 2.276$
$[8,4,2,8\rightarrow 4;2]$ $73.55$ $\bf 2.478$ $\bf 73.93$ $2.277$
int2 $[8,4,2]$ $63.35$ $\bf 3.187$ $62.29$ $\bf 2.660$
$[8,4,8\rightarrow 2]$ $62.64$ $3.289$ $62.31$ $2.670$
$[8,4,2,8\rightarrow 2]$ $62.91$ $3.138$ $\bf 62.70$ $2.673$
$[8,4,2,8\rightarrow 4;2]$ $\bf 64.32$ $3.227$ $62.60$ $2.670$